OpenStudy (anonymous):

a physician orders 0.5 oz of magnesia and alumina oral suspension antacid 4 times per day for a client with indigestion. How many milliliters of the antacid is the client to ingest in a day?

OpenStudy (anonymous):

1 oz = 29.5735296875 ml 4* 0.5 oz = 4 * 0.5 * 29.5735296875 ml = 59.147059375 ml

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

welcome